# Reclaiming “Open AI" - AI Model Serving Can Be Open Access, Yet Monetizable and Loyal

### v3

@@ Abstract
The rapid rise of AI has split model serving between open-weight distribution, which often lacks owner control and monetization, and opaque API-based approaches that risk user privacy and model transparency, forming a dichotomy that hinders an equitable AI ecosystem. This position paper introduces, rigorously formulates, and champions the Open-access, Monetizable, and Loyal (OML) paradigm for AI model serving: a foundational shift to securely distribute and serve AI models by synthesizing transparency with granular monetization and critical safety controls. We survey diverse OML constructions from theory and practice, analyze their security, performance, and practical trade-offs, outline a conceptual OML deployment protocol, and discuss market and policy implications. We assert that OML can foster a democratized, self-sustaining, and innovative AI landscape, mitigating centralized power risks. Finally, we call on the research community to further explore the broad design space of OML, spanning cryptographic, AI-native, and socio-economic mechanisms, to realize its full potential for a collaborative, accountable, and resilient AI future.
1 Introduction and Motivation - Why OML?
1.1 Era of AI
@@
Artificial Intelligence (AI) is advancing at an unprecedented pace, reshaping diverse fields from household robotics [43, 27] and superhuman game-playing [96, 97, 98] to intricate formal mathematical reasoning [23], protein structure elucidation [48, 28], accelerated drug discovery [16, 99, 94], and novel mathematical exploration [88]. The emergence of powerful generative models like GPT-4 [75, 17], OpenAI o3 [76, 44], and DeepSeek R1 [34] marks a watershed moment, heralding their potential to fundamentally transform human endeavors.
Yet, this bright promise is shadowed by a stark reality: AI’s trajectory is largely dictated by a handful of powerful organizations, whose unaccountable decisions risk unforeseen societal impacts. This concentration of power is insidiously amplified by Jevons’ Paradox [8]: As AI’s efficiency and accessibility grow, so does the demand, tightening the grip of these few entities, fostering monopolization, and stifling opportunities for a burgeoning global talent pool.
Confronting these looming threats to equitable progress, we posit an urgent imperative to democratize AI development. Our vision is an open, collaborative platform that enshrines community ownership rights, thereby unleashing broad participation in the creation, stewardship, and monetization of AI. This initiative seeks to catalyze a vibrant new era of AI commercialization, paving the way for a truly community-built open AGI.
1.2 The AI Service Landscape
@@
Modern AI development, responsible for the ubiquitous systems we see today, has been significantly shaped by open-source contributions. Until recently, core libraries and even powerful models such as BERT [25] and early iterations of GPT [75, 17] were openly available. However, as AI matured and its profound economic potential became evident, many large companies that initially embraced open development have transitioned to closed strategies. These entities have often geared their efforts towards establishing significant positions in the AI economy, with a model where others act as high-level users of the AI they build [74, 29, 57]. Consequently, AI is currently delivered to users predominantly via two distinct service models.
Concurrently, open models such as Llama [104] and DeepSeek [24] continue to improve, often remaining at par with or even surpassing closed counterparts. With virtually every industry adopting AI and the diversity of use cases expanding daily, the pool of AI innovators is growing rapidly. A significant portion of the global workforce is eager to contribute to AI development. Harnessing these collective forces could foster a more powerful, inclusive, and community-built AI ecosystem, open to widespread innovation. However, achieving this vision necessitates a new socio-technical framework that aligns incentives for AI builders, enabling them to fairly share in the economic rewards of the AI they help create, thereby fostering a more sustainable and robust development cycle.
The two dominant service models are:
@@
  • Closed with API access: In this paradigm, AI models are primarily accessed through public APIs [74, 29, 57]. For example, the OpenAI API allows users to interact with models like ChatGPT [73] and DALL-E [86]. Such centralized, closed services offer benefits like scalability and the implementation of certain safety measures, including content moderation and misuse prevention. Conversely, this approach can lead to monopolization, rent-seeking behaviors, and significant privacy concerns. Users also lack ultimate control over the paid service, as model owners can arbitrarily filter inputs, modify outputs, or change the underlying model without direct user consent. While options for fine-tuning closed models may exist, such customization is typically constrained by the limitations of the provided API. OpenAI’s GPT services and Google’s Gemini service are prime examples of this model.
@@
  • Open-weight distribution: Under this model, creators release model weights, often with their architectures, allowing users to download and run inference locally. This grants users full control over model selection, inference processes (optimizing for efficiency and privacy), and the ability to transparently build upon these models (e.g., through fine-tuning) and compose them with other AI systems. Meta’s Llama models, DeepSeek, and the wide array of models available on platforms like Hugging Face exemplify this approach. However, once models are released in this manner, creators often lose direct control over their subsequent use and face challenges in monetization. They cannot easily prevent unsafe or unethical applications—a lack of what might be termed loyalty or retained governance.
Essentially, both prevailing paradigms have substantial drawbacks. Closed AI systems force users to sacrifice control and transparency, and prominent examples like OpenAI’s ChatGPT and Google’s Gemini have faced criticism for this opacity and the risks associated with arbitrary manipulation by the provider. While the transparency and user control offered by open-weight AI are highly desirable, this model is not a panacea. It often presents significant challenges for model creators in terms of monetization, potentially undermining the incentives for innovation and leading to unsustainable production cycles. The case of DeepSeek, which relies on funding from its parent company’s hedge fund rather than its own profitability, illustrates this precariousness—an arrangement that is often impractical and fragile for broader adoption. This landscape highlights the pressing need for a more balanced approach.
1.3 New Paradigm: OML (Open-access, Monetizable, Loyal)
While both closed and open models have their merits, we aim for maximal openness, akin to today’s open-weight models, but with added monetizability and loyalty constraints. To achieve this, in this position paper, we argue that, it is both feasible and imperative that AI model serving becomes open-access, yet monetizable and loyal at the same time. We propose OML as a new AI format. Operationally, a model owner enhances their AI model (M) with a cryptographic primitive enabling these features, then openly publishes the resulting M.oml file. OML stands for Open-access, Monetizable, and Loyal.
  • Open-access. OML models are broadly accessible for local execution, similar to distributing a compiled binary rather than source code. While usable, the internal mechanisms for monetizability and loyalty remain protected. This local operation ensures the model is immutable post-distribution and offers benefits like user data privacy, consistent service quality, and on-premise advantages.
  • Monetizable. The OML framework allows model owners to derive value by requiring authorization for use, enabling them to set terms for approval. Fundamentally, every usage can be controlled and charged for, even if enforcement for unauthorized use might theoretically occur post-hoc.
  • Loyal. This property ensures the OML model inherently enforces authorization pre-hoc. The model is designed to function correctly, delivering accurate outputs, only with a valid, owner-issued permission for that specific use. This preventative control means the model will not perform as intended without prior approval, guaranteeing adherence to the owner’s policies, including ethical or safety conditions.
Monetizability and loyalty both affirm an owner’s right to govern usage, but they differ subtly. Monetizability is economic and can potentially be addressed post-hoc (e.g., via collateral for terms violations). Loyalty, however, requires pre-hoc enforcement for safe and ethical use.
1.4 Broader Vision with OML: Community-built AI
OML serves as the foundational first step towards our broader vision of a genuinely community-built AI ecosystem, which empowers every participant, from those architecting new models or meticulously fine-tuning existing ones, to those supplying crucial training data or generating insightful usage data, with a meaningful role in guiding the technology’s advancement and sharing in its success. AI builders can pioneer novel architectures and substantially enhance model capabilities, while AI users, through practical deployment and valuable data feedback, evolve into active contributors themselves, shaping the AI’s real-world efficacy.
For innovation to foster sustainable, long-term growth, it is essential to align contributor incentives with the burgeoning AI economy, ensuring that as model adoption and utility increase, all who created or improved them share in the rewards. Figure 1 clearly contrasts the limitations of current one-way platforms, which often lack vital feedback mechanisms, against our proposed closed-loop system. In this dynamic system, active user engagement not only continuously refines the platform but also fosters a virtuous cycle by aligning incentives for sustained, collaborative innovation.
With OML as the cornerstone, we aim to truly democratize AI, cultivating an open and vibrant ecosystem that actively empowers both contributors and users, distinct from the restrictive nature of de facto closed models like “OpenAI”. Our strategic approach with OML is designed to ensure that widespread community participation and a flourishing culture of innovation can coexist harmoniously with the robust commercial viability and monetization of the developed models.
Figure 1: On the left, a one-way interaction is depicted, where AI platforms deliver outputs to users without feedback. On the right, the process evolves into a closed loop, where user feedback influences the AI platform, aligning incentives and enabling continuous improvement in an open environment.
2 OML Formulation: Open-access, Monetizable, and Loyal Model Serving
In this section, we introduce the Open-access, Monetizable, and Loyal (OML) model serving paradigm. We begin by introducing the OML protocol, outlining its idealized workflow, core design components, and associated notations (Section 2.1). Then, we delineate the fundamental properties that an OML-formatted AI model, denoted $M_{oml}$, must satisfy (Section 2.2). We further discuss the adversary model and the security guarantees OML primitives aim to provide against various attack vectors, formalized through rigorous security definitions (Section 2.3). This framework is the foundation of a detailed exploration of potential canonical approaches to achieving OML and follow-up research, which will be covered in Section 3.
2.1 The OML Protocol and Design Space
Definition 2.1 (OML-formatted Model).
An OML-formatted model, denoted $M_{oml}$, is derived from an original plain-text model $M$ through a transformation function $OMLize(M)\rightarrow M_{oml}$. This transformation embeds mechanisms within $M_{oml}$ designed to ensure it exhibits the core properties of open-access, monetizability, and loyalty, as formally defined in Section 2.2.
Ideal OML Workflow. Let $M:\mathcal{X}\to\mathcal{Y}$ be the original plain-text model. An OML-formatted version, $M_{oml}$, is constructed. Let $k_{own}$ be $\mathcal{O}$’s secret ownership key. The protocol involves an input transformation function $h:\mathcal{X}\to\mathcal{H}$ (e.g., hashing, or extracting identifying features from $x$), and a permission generation function $\sigma:\mathcal{H}\times\mathcal{K}_{own}\to\mathcal{P}$, where $\mathcal{P}$ is the space of permission tokens. For a given input $x$, the permission token is $p_{x}=\sigma(h(x),k_{own})$. The OML model $M_{oml}$ takes as input a pair $s_{x}=(x,p_{x})\in\mathcal{X}\times\mathcal{P}$. Let $\mathcal{S}=\mathcal{X}\times\mathcal{P}$ be the input space for $M_{oml}$.
The usage of $M_{oml}$ is as follows:
@@ 2.1 The OML Protocol and Design Space
    Authorization: To process an input $x$, an authorized user obtains $p_{x}=\sigma(h(x),k_{own})$ from an owner-controlled mechanism (e.g., a platform $\Pi_{\mathcal{O}}$ that, upon receiving $h(x)$ and confirming authorization, returns $p_{x}$). The user then queries $M_{oml}$ with $s_{x}=(x,p_{x})$.
@@
    Correctness Guarantee: If $p_{x}$ is valid, $M_{oml}(s_{x})$ produces an output $y_{oml}\in\mathcal{Y}$. Ideally, $y_{oml}$ should be close to the output of the original model $M(x)$. Formally, for a distance metric $d:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$, $d(M_{oml}(x,p_{x}),M(x))\leq\epsilon_{utility}$ for some small $\epsilon_{utility}\geq 0$.
@@
    Monetization Guarantee: If an adversary $\mathcal{A}$ attempts to query $M_{oml}$ with an invalid input $s^{\prime}_{x}=(x,p^{\prime}_{x})$ where $p^{\prime}_{x}$ is not a valid token for $x$ (e.g., $p^{\prime}_{x}$ is guessed, replayed from another input, or $s^{\prime}_{x}=x$ if $M_{oml}$ expects a pair), then $M_{oml}(s^{\prime}_{x})$ should yield an inaccurate or useless output $y^{\prime}_{oml}$. Formally, $d(M_{oml}(s^{\prime}_{x}),M(x))>\epsilon_{robust}$ for some $\epsilon_{robust}>\epsilon_{utility}$ when $p^{\prime}_{x}$ is far from any valid token, i.e. $d(s^{\prime}_{x},s_{x})>k\cdot\epsilon_{robust}$ for any valid $s_{x}=(x,p_{x})$.
@@
    Efficiency: The computational overhead of using $M_{oml}$ should be minimal. Let $T(F,z)$ be the computation overhead to compute $F(z)$. Formally, we require $T(M_{oml},(x,p_{x}))\leq(1+\epsilon_{overhead})T(M,x)$ for a small $\epsilon_{overhead}\geq 0$.
This workflow (illustrated in Figure 2) aims to protect ownership without substantially degrading model performance or efficiency for legitimate, authorized use, and the quality of an OMLization process can be characterized by the parameter set $(\epsilon_{utility},\epsilon_{robust},\epsilon_{overhead},k)$.
Figure 2: An illustration of the ideal OML protocol. The system on the LHS is the original model, $M:\mathcal{X}\to\mathcal{Y}$. The system on the RHS is the OMLized model, $M_{oml}:\mathcal{S}\to\mathcal{Y}$. To produce an output $y\approx M(x)$, the input to $M_{oml}$ must be $s_{x}=(x,p_{x})$, where $p_{x}$ is a valid permission for $x$.
Key Components of the OML Design Space. The realization of an OML scheme involves careful design of:

### v4

# OML: A Primitive for Reconciling Open Access with Owner Control in AI Model Distribution

@@ Abstract
The current paradigm of AI model distribution presents a fundamental dichotomy: models are either closed and API-gated, sacrificing transparency and local execution, or openly distributed, sacrificing monetization and control. We introduce OML (Open-access, Monetizable, and Loyal AI Model Serving), a primitive that enables a new distribution paradigm where models can be freely distributed for local execution while maintaining cryptographically enforced usage authorization. We are the first to introduce and formalize this problem, introducing rigorous security definitions tailored to the unique challenge of white-box model protection: model extraction resistance and permission forgery resistance. We prove fundamental bounds on the achievability of OML properties and characterize the complete design space of potential constructions, from obfuscation-based approaches to cryptographic solutions. To demonstrate practical feasibility, we present OML 1.0, a novel OML construction leveraging AI-native model fingerprinting coupled with crypto-economic enforcement mechanisms. Through extensive theoretical analysis and empirical evaluation, we establish OML as a foundational primitive necessary for sustainable AI ecosystems. This work opens a new research direction at the intersection of cryptography, machine learning, and mechanism design, with critical implications for the future of AI distribution and governance.
1 Introduction
@@
Artificial Intelligence (AI) is advancing at an incredible pace, reshaping diverse fields from household robotics [39, 23] and superhuman game-playing [91, 92, 93] to intricate formal mathematical reasoning [20], protein structure elucidation [44, 24], accelerated drug discovery [15, 94, 88], and novel mathematical exploration [82]. The emergence of powerful generative models like GPT series [70, 16], OpenAI o3 [71, 40], and DeepSeek R1 [30] marks a watershed moment, heralding their potential to fundamentally transform human endeavors.
Yet, the rapid advancement of artificial intelligence has created an unprecedented challenge in model distribution. Current approaches force an unnecessary tradeoff between accessibility and control, limiting both innovation and sustainable development of AI systems. This paper introduces and formalizes OML, a primitive that reconciles these conflicting requirements.
1.1 The Fundamental Distribution Problem
@@
Modern AI development, responsible for the ubiquitous systems we see today, has been significantly shaped by open-source contributions. Until recently, core libraries and powerful models such as BERT [21] and early iterations of GPT [70, 16] were openly available. However, as AI matured and its profound economic potential became evident, many large companies that initially embraced open development have transitioned to closed strategies. These entities have often geared their efforts towards establishing significant positions in the AI economy, with a model where others act as high-level users of the AI they build [69, 25, 53]. Consequently, AI is currently delivered to users predominantly via the following two dominant paradigms, each with critical limitations.
@@
Closed API Services: In this paradigm, AI models are primarily accessed through public APIs [69, 25, 53]. Platforms like OpenAI’s GPT and Anthropic’s Claude maintain complete control over model execution, enabling monetization and usage governance. Such centralized, closed services offer benefits like scalability and the implementation of certain safety measures, including content moderation and misuse prevention. Conversely, this approach can lead to monopolization, rent-seeking behaviors, and significant privacy concerns. Users also lack ultimate control over the paid service, as model owners can arbitrarily filter inputs, modify outputs, or change the underlying model without direct user consent. While options for fine-tuning closed models may exist, such customization is typically constrained by the limitations of the provided API. Users cannot verify model behavior, ensure data privacy, or maintain operational independence.
@@
Open-weight distribution: In this paradigm, creators release model weights, often with their architectures, allowing users to download and run inference locally. Platforms like Hugging Face enable unrestricted model distribution, providing transparency, local execution, and modification capabilities. This grants users full control over model selection, inference processes, and the ability to build upon these models (e.g., through fine-tuning) and compose them with other AI systems. Meta’s Llama model series [99], DeepSeek [30], and the wide array of models available on platforms like Hugging Face exemplify this approach. However, once models are released in this manner, creators lose direct control over their subsequent use, cannot enforce safety constraints and prevent unethical applications, and lack mechanisms for sustainable monetization. This creates a tragedy of the commons where innovation is undersupported.
This dichotomy is not merely a business model choice but represents a fundamental technical limitation in our current infrastructure. We lack the primitives necessary to enable models that are simultaneously open-access for local execution and controlled for authorized and ethical usage.
1.2 OML: Technical Requirements for Ideal Resolution
While both closed and open models offer distinct advantages, we pursue maximal openness comparable to current open-weight distributions, augmented with mechanisms for owner control. An ideal solution must therefore reconcile three seemingly contradictory properties, which we term “OML":
  • [O] Open-access. Models must be freely distributable for local execution, analogous to compiled binaries where functionality is accessible while implementation details remain protected. Once distributed, models become immutable artifacts independent of their creators, enabling user data privacy, consistent service quality, and on-premise deployment.
  • [M] Monetizable. The framework enables model owners to capture economic value through granular, per-inference authorization mechanisms. Each model invocation requires individual permission from model owners, and an economic Nash Equilibrium, where rational users obtain proper authorization (such as purchasing access tokens) rather than attempting circumvention, guarantees that the all agreements and licenses are enforced.
  • [L] Loyal. Models must technically enforce owner-defined policies through pre-hoc authorization verification. The model produces high-utility outputs exclusively when presented with valid, cryptographically-bound permissions, ensuring compliance with safety and ethical constraints before computation rather than through legal remedies after violation.
Note that, while monetizability and loyalty both enable governance, they address fundamentally different requirements. Monetizability concerns economic sustainability and can leverage post-hoc mechanisms such as collateral-backed compliance. Loyalty demands pre-hoc technical enforcement to prevent generation of harmful or policy-violating outputs before they occur.
This challenge becomes particularly acute under white-box access conditions where users possess complete visibility into model weights, architecture, and computation flow. Unlike traditional software where obfuscation can leverage discrete execution paths and control flow complexity, neural networks present continuous, differentiable computations that resist conventional protection mechanisms. Any solution must therefore leverage properties unique to machine learning systems rather than adapting existing software protection paradigms.
1.3 Our Contributions
Core contribution. We are the first to identify and formalize the OML challenge, a primitive that enables AI models to be distributed openly while maintaining cryptographically enforced usage control under white-box access. This paper establishes the foundations for OML through:
@@ 1.3 Our Contributions
[…]
@@
    Security Framework: We establish novel security games for white-box model protection (model extraction resistance, permission forgery resistance) where adversaries have complete visibility into weights and computation, establishing security standards for OML realization.
@@
    Design Space Characterization: We analyze a wide array of potential approaches (obfuscation, TEEs, cryptography) with their fundamental tradeoffs and theoretical bounds.
@@
    Feasibility Demonstration: We introduce and instantiate OML 1.0, a practical construction using AI-native fingerprinting that achieves “next-day security" with negligible $\epsilon_{overhead}$. Experiments and empirical validation further demonstrates feasibility of our approach.
Our primary contribution is establishing OML as a well-defined primitive and demonstrating its feasibility, thereby opening a new research direction at the intersection of cryptography, machine learning, and mechanism design. The challenge of protecting neural networks under white-box access while preserving utility represents one of the most difficult problems in cryptography and AI, requiring fundamentally new techniques beyond classical approaches. We identify critical open problems spanning theoretical questions (tight complexity bounds, connections to program obfuscation, optimal constructions) and practical challenges (robustness under fine-tuning, compositional security, efficient real-time enforcement).
The full realization of OML will likely require years of sustained research effort from the community, comparable to the decades-long development of practical homomorphic encryption or secure multiparty computation. By providing the first formal framework and proving initial feasibility through OML 1.0, we aim to catalyze the long-term research program necessary to advance OML from primitive concept to robust, deployable infrastructure for AI distribution.
1.4 Significance and Impact
Through the OML primitive, we can address critical safety and privacy challenges faced in AI development today, while enabling a more collaborative ecosystem for AI development tomorrow:
Privacy-Preserving Local Execution: OML enables organizations handling sensitive data, e.g. healthcare providers, financial institutions, government agencies, to deploy state-of-the-art models without exposing confidential information to external APIs. Medical institutions can run diagnostic models on patient data locally while still compensating model creators. Financial firms can apply fraud detection models without sharing transaction patterns. This solves the current impossibility where organizations must choose between using inferior models or violating privacy requirements.
Technical Enforcement of Ethical and Legal Usage: By requiring pre-hoc cryptographic authorization for each inference, OML provides robust mechanisms for enforcing safety policies and any signed user agreements on open-weight models that currently rely solely on terms of service today.
Enabling a More Colloborative Ecosystem for AI: Most fundamentally, OML creates the technical infrastructure for distributed AI development where every contribution, from initial training to incremental improvements, can be tracked and rewarded. Figure 1 illustrates how OML transforms the current one-way distribution into a sustainable ecosystem where usage generates returns for all contributors, creating economic incentives for continuous improvement rather than one-time model releases. Collaboration with positive feedback loop and proper incentives, rather than monopolization, guarantees that AI development is on the right track that benefits the humanity.
Figure 1: OML enables transition from one-way model distribution to bidirectional value flow.
Left: Current paradigm where models are distributed without feedback or compensation mechanisms.
Right: OML-enabled ecosystem where usage generates returns for all contributors, incentivizing continuous collaborative improvement.
These three impacts are mutually reinforcing: privacy-preserving local execution expands the market for AI models; license enforcement makes broader deployment responsible; and collaborative development efforts ensure sustainable innovation. Together, they address the fundamental limitations preventing AI from achieving its potential as a broadly beneficial technology.
2 The OML Primitive: Formal Definition
The transformation of AI models from static artifacts to dynamic, controllable assets requires a fundamental reconceptualization of how we distribute and govern computational intelligence. In this section, we formalize the Open-access, Monetizable, and Loyal (OML) primitive, a framework that enables white-box model distribution while preserving ownership rights and enforcing usage policies.
2.1 Properties and Design Space
Consider an AI model $M$ as an economic asset: valuable, replicable, and vulnerable to unauthorized extraction once distributed. The OML primitive transforms this vulnerable asset into a controlled artifact that retains its utility while coupling high-quality outputs to owner authorization.
We begin by establishing our notation framework, which we will use throughout this section.
Table 1: Notation and Core Components of the OML Framework
Symbol Description
$M:\mathcal{X}\!\to\!\mathcal{Y}$ Original model mapping inputs to outputs
$M_{\text{oml}}$ OML-formatted model with embedded authorization
$h:\mathcal{X}\!\to\!\mathcal{H}$ Input-binding transform (e.g., cryptographic commitment)
$\sigma:\mathcal{H}\!\times\!\mathcal{K}_{\text{own}}\!\to\!\mathcal{P}$ Permission token generator
$k_{\text{own}}$ Owner’s secret key; $vk_{\text{own}}$ denotes optional public verifier
$p_{x}=\sigma(h(x),k_{\text{own}})$ Permission token cryptographically bound to input $x$
$d(\cdot,\cdot)$ Task-appropriate distance or divergence metric
$\epsilon_{\text{utility}}$ Maximum fidelity loss on authorized queries
$\epsilon_{\text{robust}}$ Minimum degradation on unauthorized queries
$\epsilon_{\text{overhead}}$ Relative computational overhead bound
With this notation established, we can now formally define the OML transformation, a process that embeds authorization logic so deeply within the model’s computational graph that removing it becomes computationally equivalent to retraining the model from scratch.
Definition 2.1 (OMLized Model).
Given an original model $M:\mathcal{X}\!\to\!\mathcal{Y}$, an OMLization process
$\mathrm{OMLize}(M;h,\sigma,\mathrm{params})\ \longrightarrow\ M_{oml},$
produces a locally executable artifact that operates on input-token pairs $(x,p)$. For each input $x\in\mathcal{X}$, authorization requires a valid token $p_{x}=\sigma(h(x),k_{own})$ computed with owner’s secret key $k_{own}\in\mathcal{K}_{own}$. Informally, $M_{oml}$ behaves as $M$ on authorized inputs and degrades otherwise.
This definition captures the essence of controlled distribution: the model remains functionally accessible but computationally gated. To understand how this works in practice, we present the idealized OML workflow, which demonstrates how authorization, utility preservation, and security enforcement interact.
Promises of an Ideal OML.
Let $d:\mathcal{Y}\!\times\!\mathcal{Y}\!\to\!\mathbb{R}_{\geq 0}$ denote a distance metric and $T(F,z)$ the computational cost of evaluating function $F$ at input $z$. A correct OMLization satisfies: